State the relationship between the volume of a given mass of gas, and its pressure at constant temperature
1
Expert's answer
2019-11-18T06:28:26-0500

Solution.

Since there is an isothermal process (Boyle-Marriott Law), we write the equation of state of the ideal gas:

"p1 \times V1 = p2 \times V2"

The relationship between the volume of a given mass of gas and pressure is as follows:

"p \times V = const"

Answer:

"p \times V = const"